Please help me in solving this question.
Vijay is x years old and his father is nine times the square of vijay's age. Thirty-two years hence,his father's age will be double of his age then. Find their present ages.

Answer by Alan3354(69443)

Vijay is x years old and his father is nine times the square of vijay's age. Thirty-two years hence, his father's age will be double of his age then. Find their present ages.
--------
F = 9x^2
F+32 = 2*(x+32) = 2x+64 --> F = 2x+32
---
9x^2 = 2x + 32
9x^2 - 2x - 32 = 0
(x-2)*(9x + 16) = 0
x = 2, -16/9 *** Ignore the negative solution.
-------
x = 2
F = 36
=====================
Thirty-two years hence, his father's age will be double of his age then.
Not clear when "then" is.